OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0. a. x = -1 and x = -5 b. 3 plus or minus square root of 2 c. negative 3 plus or minus square root of 2 d. quantity of negative 3 plus or minus square root of 2 all over 2 Solve 3x2 + 4x = 2. a. quantity of 2 plus or minus square root of 10 all over 6 b. quantity of negative 2 plus or minus 2 square root of 10 all over 3 c. quantity of negative 2 plus or minus square root of 10 all over 3 d. quantity of 4 plus or minus square root of 10 all over 3 Solve 3x2 = -30x + 24. a. quantity of negative 10 plus or minus square root of 33 all over 2

OpenStudy (anonymous):

easy stuff bro

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

c

OpenStudy (anonymous):

\[x^2+6x+7=0\] \[x^2+6x=-7\] \[(x+3)^2=-7+3^2=-7+9=2\] \[x+3=\pm\sqrt{2}\] \[x=-3\pm\sqrt{2}\] is the easiest way

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

mutual medals ^_^

OpenStudy (anonymous):

gah im having a a hard time with these.. what are the other two -_-

OpenStudy (anonymous):

the other one is

OpenStudy (anonymous):

SR of 2/3

OpenStudy (anonymous):

+ or - SR

OpenStudy (anonymous):

okay? which letters?

OpenStudy (anonymous):

medal pl0x

OpenStudy (anonymous):

Solve 3x2 + 4x = 2. a. quantity of 2 plus or minus square root of 10 all over 6 b. quantity of negative 2 plus or minus 2 square root of 10 all over 3 c. quantity of negative 2 plus or minus square root of 10 all over 3 d. quantity of 4 plus or minus square root of 10 all over 3 Solve 3x2 = -30x + 24. a. quantity of negative 10 plus or minus square root of 33 all over 2 b. negative 5 plus or minus 2 square root of 33 c. negative 10 plus or minus square root of 33 d. negative 5 plus or minus square root of 33

OpenStudy (anonymous):

i'll give you the answers for a medal of honour

OpenStudy (anonymous):

deal

OpenStudy (anonymous):

medal of honour plox

OpenStudy (anonymous):

okay